# Existence and non-existence of consistent estimators in supercritical controlled branching processes

### v1

@@ 5 Proofs
  3. (iii)
    $||\mathcal{L}_{\tilde{\xi}_{Z}(z)}-\mathcal{L}_{(\tilde{\xi}_{Z}(z)+1)}||_{TV}$
    $=1-\sum_{j=0}^{\infty}\mathbbm{P}\bigg{(}\sum_{i=1}^{\tilde{\phi}_{Z}(z)}\xi_{Z,i}=j\bigg{)}\wedge\mathbbm{P}\bigg{(}\sum_{i=1}^{\tilde{\phi}_{Z}(z)}\xi_{Z,i}=j-1\bigg{)}$
    $\leq 1-\mathbbm{P}\big{(}\tilde{\phi}_{Z}(z)\geq 1\big{)}\cdot\sum_{k=0}^{\infty}\Bigg{(}\mathbbm{P}(\xi_{Z,1}=x)\cdot\mathbbm{P}\bigg{(}\sum_{i=2}^{\tilde{\phi}_{Z}(z)}\xi_{Z,i}=k\bigg{)}\wedge\mathbbm{P}(\xi_{Z,1}=x-1)\cdot\mathbbm{P}\bigg{(}\sum_{i=2}^{\tilde{\phi}_{Z}(z)}\xi_{Z,i}=k\bigg{)}\Bigg{)}$
    $=1-\mathbbm{P}\big{(}\tilde{\phi}_{Z}(z)\geq 1\big{)}\big{(}\mathbbm{P}(\xi_{Z,1}=x)\wedge\mathbbm{P}(\xi_{Z,1}=x-1)\big{)}.$
    which we can solve to see that $\mathbbm{P}\big{(}\tilde{\phi}_{Z}(z)\geq 1\big{)}>t_{Z}^{2}/(t_{Z}^{2}+\lambda_{Z}^{3}m_{Z}^{2}b_{Z})>0$ for all $z\geq M_{3}$. Hence there exists a positive constant $\eta_{Z}$ such that, for all $z\geq M_{3}$, $||\mathcal{L}_{\tilde{\xi}_{Z}(z)}-\mathcal{L}_{(\tilde{\xi}_{Z}(z)+1)}||_{TV}\leq\eta_{Z}<1$. Further, for $\eta^{\prime}_{Z}:=\eta_{Z}\vee 3/4$, we obtain $1+4(l_{Z}(z)-1)(1-\eta_{Z})\geq 4(1-\eta^{\prime}_{Z})l_{Z}(z)$.
Inputting these results into Lemma 5.5, for $W_{Z}(z)\sim\text{DN}\big{(}l_{Z}(z)\cdot\tilde{m}_{Z}(z),\;l_{Z}(z)\cdot\tilde{\sigma}_{Z}^{2}(z)\big{)}$ and $z\geq M_{3}$, we obtain
$||\mathcal{L}_{Z_{1}|Z_{0}=z}-\mathcal{L}_{W_{Z}(z)}||_{TV}$
$\leq\sqrt{\frac{2}{\pi}}\left(\frac{3\tilde{\rho}_{Z}(z)}{\tilde{\sigma}_{Z}^{2}(z)}+2\right)\Big{(}1+4(l_{Z}(z)-1)\left(1-||\mathcal{L}_{\tilde{\xi}_{Z}(z)}-\mathcal{L}_{(\tilde{\xi}_{Z}(z)+1)}||_{TV}\right)\Big{)}^{-\frac{1}{2}}$
and hence see that
$||\mathcal{L}_{Z_{1}|Z_{0}=z}-\mathcal{L}_{W_{Z}(z)}||_{TV}=O(z^{-1/2}).$
Repeating the same process for $\{X_{n},z_{0}\}$, we can form an analogous bound on $||\mathcal{L}_{X_{1}|X_{0}=z}-\mathcal{L}_{W_{X}(z)}||_{TV}$, for $W_{X}(z)\sim\text{DN}\big{(}l_{X}(z)\cdot\tilde{m}_{X}(z),\;l_{X}(z)\cdot\tilde{\sigma}_{X}^{2}(z)\big{)}$, with
$||\mathcal{L}_{X_{1}|X_{0}=z}-\mathcal{L}_{W_{X}(z)}||_{TV}=O(z^{-1/2}).$
It follows from the triangle inequality that
$||\mathcal{L}_{Z_{1}|Z_{0}=z}-\mathcal{L}_{X_{1}|X_{0}=z}||_{TV}\leq||\mathcal{L}_{Z_{1}|Z_{0}=z}-\mathcal{L}_{W_{Z}(z)}||_{TV}+||\mathcal{L}_{W_{Z}(z)}-\mathcal{L}_{W_{X}(z)}||_{TV}+||\mathcal{L}_{X_{1}|X_{0}=z}-\mathcal{L}_{W_{X}(z)}||_{TV},$
so the result will follow if we can show that there exists $s>0$ such that $||\mathcal{L}_{W_{Z}(z)}-\mathcal{L}_{W_{X}(z)}||_{TV}=O(z^{-s})$.
Lemma 5.7 allows us to bound
$||\mathcal{L}_{W_{Z}(z)}-\mathcal{L}_{W_{X}(z)}||_{TV}$
$=\frac{3|\text{Var}(Z_{1}|Z_{0}=z)-\text{Var}(X_{1}|X_{0}=z)|}{2(l_{Z}(z)\cdot\tilde{\sigma}_{Z}^{2}(z))}+\frac{|\mathbbm{E}(Z_{1}|Z_{0}=z)-\mathbbm{E}(X_{1}|X_{0}=z)|}{2\sqrt{l_{Z}(z)\cdot\tilde{\sigma}_{Z}^{2}(z)}}$
for $z\geq M_{3}$, and since there exists $r<1$ such that $|\mathbbm{E}(Z_{1}|Z_{0}=z)-\mathbbm{E}(X_{1}|X_{0}=z)|=O(z^{r/2})$ and $|\text{Var}(Z_{1}|Z_{0}=z)-\text{Var}(X_{1}|X_{0}=z)|=O(z^{r})$, we see that $||\mathcal{L}_{W_{Z}(z)}-\mathcal{L}_{W_{X}(z)}||_{TV}=O(z^{(r-1)/2})$, so that $||\mathcal{L}_{W_{Z}(z)}-\mathcal{L}_{W_{X}(z)}||_{TV}=O(z^{-s})$ for $s=(1-r)/2$. $\blacksquare$
We omit the proof of Theorem 3.10, which can be shown using the same argument as Theorem 3.5, proved above.
5.4 Proofs for Section 3.4
To prove Lemma 3.11, we first introduce the following lemma:
Lemma 5.8.
Let $\phi_{Z}(z)$ and $\phi_{X}(z)$ be two control functions satisfying (8) and with $\liminf_{z\to\infty}\nu^{2}(z)/z>0$, linearly-divisible such that (13) is satisfied. Then, if there exists $r<1$ such that
then there exists $q>0$ such that
$||\mathcal{L}_{\phi_{Z}(z)}-\mathcal{L}_{\phi_{X}(z)}||_{TV}=O(z^{-q}).$
Proof. Let $\phi_{Z}(z)$ and $\phi_{X}(z)$ be as stated in the lemma. The triangle inequality allows us to form the bound
$||\mathcal{L}_{\phi_{Z}(z)}-\mathcal{L}_{\phi_{X}(z)}||_{TV}\leq||\mathcal{L}_{\phi_{Z}(z)}-\mathcal{L}_{W_{\phi_{Z}}(z)}||_{TV}+||\mathcal{L}_{W_{\phi_{Z}}(z)}-\mathcal{L}_{W_{\phi_{X}}(z)}||_{TV}+||\mathcal{L}_{\phi_{X}(z)}-\mathcal{L}_{W_{\phi_{X}}(z)}||_{TV},$
Since $\phi_{Z}(\cdot)$ is linearly-divisible, for a given $z\in\mathbbm{N}_{0}$, we can decompose $\phi_{Z}(z)\stackrel{{\scriptstyle d}}{{=}}\sum_{i=1}^{l_{Z}(z)}\tilde{\phi}_{Z,i}(z)$, where $\{\tilde{\phi}_{Z,i}(z)\}_{i\in\mathbbm{N}_{1}}$ are all i.i.d. We define $\tilde{\varepsilon}_{Z}(z):=\mathbbm{E}\tilde{\phi}_{Z}(z)$, $\tilde{\nu}^{2}_{Z}(z):=\text{Var}(\tilde{\phi}_{Z}(z))$, and $\tilde{\upsilon}_{Z}(z):=\mathbbm{E}(\tilde{\phi}_{Z}(z)-\mathbbm{E}\tilde{\phi}_{Z}(z))^{3}$, so that
$\tilde{\varepsilon}_{Z}(z)=\frac{\varepsilon_{Z}(z)}{z}\cdot\frac{z}{l_{Z}(z)},\quad\tilde{\nu}^{2}_{Z}(z)=\frac{\nu^{2}_{Z}(z)}{z}\cdot\frac{z}{l_{Z}(z)},\quad\text{and}\quad\tilde{\upsilon}_{Z}(z)=\frac{\upsilon_{Z}(z)}{z}\cdot\frac{z}{l_{Z}(z)}.$
$\tilde{\varepsilon}_{Z}(z)\leq a_{Z}\lambda_{Z},\quad\tilde{\nu}_{Z}^{2}(z)\leq b_{Z}\lambda_{Z},\quad\text{and}\quad\tilde{\upsilon}_{Z}(z)\leq c_{Z}\lambda_{Z}$ (26)
for $z\geq M_{1}$. We can then note that
@@
  3. (iii)
    Because the $\tilde{\phi}_{Z,i}(z)$’s satisfy (13), we have $||\mathcal{L}_{\tilde{\phi}_{Z}(z)}-\mathcal{L}_{(\tilde{\phi}_{Z}(z)+1)}||_{TV}\leq 1-\eta_{Z}$.
Then, by Lemma 5.5,
$||\mathcal{L}_{\phi_{Z}(z)}-\mathcal{L}_{W_{\phi_{Z}}(z)}||_{TV}$ $\leq\sqrt{\frac{2}{\pi}}\left(\frac{3\lambda_{Z}R_{Z}}{t_{Z}}+2\right)\frac{1}{\sqrt{1+4\eta_{Z}(z/\lambda_{Z}-1)}}+\left(5+3\sqrt{\frac{\pi}{8}}\right)\frac{\sqrt{\lambda_{Z}^{3}}R_{Z}}{\sqrt{t_{Z}^{3}\cdot z}}+\frac{\sqrt{\lambda_{Z}}}{2\sqrt{2\pi t_{Z}\cdot z}}$
for $z\geq M_{3}$, so that we see
$||\mathcal{L}_{\phi_{Z}(z)}-\mathcal{L}_{W_{\phi_{Z}}(z)}||_{TV}=O(z^{-1/2}).$
Analogously, we can find
$||\mathcal{L}_{\phi_{X}(z)}-\mathcal{L}_{W_{\phi_{X}}(z)}||_{TV}=O(z^{-1/2}).$
It remains to bound $||\mathcal{L}_{W_{\phi_{Z}}(z)}-\mathcal{L}_{W_{\phi_{X}}(z)}||_{TV}$. By Lemma 5.7, and since $\nu^{2}_{Z}(z)>t_{Z}z$ for all $z\geq M_{2}$, we see that
$||\mathcal{L}_{W_{\phi_{Z}}(z)}-\mathcal{L}_{W_{\phi_{X}}(z)}||_{TV}$ $\leq\frac{3|\nu^{2}_{Z}(z)-\nu^{2}_{X}(z)|}{2t_{Z}\cdot z}+\frac{|\varepsilon_{Z}(z)-\varepsilon_{X}(z)|}{2\sqrt{t_{Z}\cdot z}}$
for $z\geq M_{2}$. Since $|\varepsilon_{Z}(z)-\varepsilon_{X}(z)|=O(z^{r/2})$ and $|\nu^{2}_{Z}(z)-\nu^{2}_{X}(z)|=O(z^{r})$, we see that $||\mathcal{L}_{W_{\phi_{Z}}(z)}-\mathcal{L}_{W_{\phi_{X}}(z)}||_{TV}=O(z^{(r-1)/2})$. Hence $||\mathcal{L}_{\phi_{Z}(z)}-\mathcal{L}_{\phi_{X}(z)}||_{TV}=O(z^{-q})$, for $q:=(1-r)/2$. $\blacksquare$
Proof of Lemma 3.11. Let $\{Z_{n},z_{0}\}$ and $\{X_{n},z_{0}\}$ be two supercritical CBPs with control functions satisfying (8) and with $\liminf_{z\to\infty}\nu^{2}(z)/z>0$, that are linearly-divisible into random variables satisfying (13), and with offspring distributions having finite third moments and lattice size one. Assume that $m_{Z}=m_{X}$ and $\sigma^{2}_{Z}=\sigma^{2}_{X}$, and there exists $r<1$ such that $|\varepsilon_{Z}(z)-\varepsilon_{X}(z)|=O(z^{r/2})$ and $|\nu^{2}_{Z}(z)-\nu^{2}_{X}(z)|=O(z^{r})$.
Since $m_{Z}=m_{X}$, $\sigma^{2}_{Z}=\sigma^{2}_{X}$, and $\xi_{Z}$ and $\xi_{X}$ both have finite third moments and lattice size one, it follows from [26, Theorem 9] that, for a given $u\in\mathbbm{N}_{1}$, there exists a constant $c$ depending on $\xi_{Z}$ and $\xi_{X}$ such that
$||\mathcal{L}_{Z_{1}|\phi_{Z}(Z_{0})=u}-\mathcal{L}_{X_{1}|\phi_{X}(X_{0})=u}||_{TV}=||\mathcal{L}_{\sum_{i=1}^{u}\xi_{Z,i}}-\mathcal{L}_{\sum_{i=1}^{u}\xi_{X,i}}||_{TV}\leq\frac{c}{\sqrt{u}}.$
$||\mathcal{L}_{(\phi_{Z}(Z_{0}),\,Z_{1})|Z_{0}=z_{0}}-\mathcal{L}_{(\phi_{X}(X_{0}),\,X_{1})|X_{0}=z_{0}}||_{TV}\leq||\mathcal{L}_{\phi_{Z}(z_{0})}-\mathcal{L}_{\phi_{X}(z_{0})}||_{TV}+\mathbbm{P}(\phi_{Z}(z_{0})\leq N)+\frac{c}{\sqrt{N+1}}.$
$||\mathcal{L}_{(\phi_{Z}(Z_{0}),\,Z_{1})|Z_{0}=z_{0}}-\mathcal{L}_{(\phi_{X}(X_{0}),\,X_{1})|X_{0}=z_{0}}||_{TV}$ $\leq||\mathcal{L}_{\phi_{Z}(z_{0})}-\mathcal{L}_{\phi_{X}(z_{0})}||_{TV}+\frac{\nu^{2}_{Z}(z_{0})}{\big{(}\varepsilon_{Z}(z_{0})-\lfloor\alpha\cdot\varepsilon_{Z}(z_{0})\rfloor\big{)}^{2}}$
$\quad+\frac{c}{\sqrt{\lfloor\alpha\cdot\varepsilon_{Z}(z)\rfloor+1}}$
$\leq||\mathcal{L}_{\phi_{Z}(z_{0})}-\mathcal{L}_{\phi_{X}(z_{0})}||_{TV}+\frac{\nu^{2}_{Z}(z_{0})}{(1-\alpha)^{2}\varepsilon_{Z}^{2}(z_{0})}+\frac{c}{\sqrt{\alpha\cdot\varepsilon_{Z}(z_{0})}}.$
$||\mathcal{L}_{(\phi_{Z}(Z_{0}),\,Z_{1})|Z_{0}=z_{0}}-\mathcal{L}_{(\phi_{X}(X_{0}),\,X_{1})|X_{0}=z_{0}}||_{TV}\leq||\mathcal{L}_{\phi_{Z}(z_{0})}-\mathcal{L}_{\phi_{X}(z_{0})}||_{TV}+\frac{b}{(1-\alpha)^{2}m_{Z}^{2}\cdot z_{0}}+\frac{c}{\sqrt{\alpha m_{Z}\cdot z_{0}}}.\$
From Lemma 5.8, we know that there exists $\tilde{q}>0$ such that $||\mathcal{L}_{\phi_{Z}(z_{0})}-\mathcal{L}_{\phi_{X}(z_{0})}||_{TV}=O(z_{0}^{-\tilde{q}})$, so that, for $q:=\tilde{q}\wedge 1/2$,
$||\mathcal{L}_{(\phi_{Z}(Z_{0}),\,Z_{1})|Z_{0}=z_{0}}-\mathcal{L}_{(\phi_{X}(X_{0}),\,X_{1})|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q}).$
$\blacksquare$
We omit the proof of Theorem 3.13, which can again be shown using the same argument as Theorem 3.5, proved above.

### v2

@@ 6 Proofs of consistency results
    We can decompose
    $\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\phi(Z_{k-1})=\frac{(Z_{n}-m\cdot\phi(Z_{n-1}))^{2}}{n\cdot\phi(Z_{n-1})}\cdot\sum_{k=1}^{n}\frac{\phi(Z_{k-1})}{\phi(Z_{n-1})}.$
    and have already shown in Lemma 6.5 that $\sum_{k=1}^{n}\frac{\phi(Z_{k-1})}{\phi(Z_{n-1})}$ is asymptotically bounded in probability by a constant. Therefore the result $\frac{(m-\tilde{m}_{n})^{2}}{n}\sum_{k=1}^{n}\phi(Z_{k-1})\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$ will follow if we show that $\frac{(Z_{n}-m\cdot\phi(Z_{n-1}))^{2}}{n\cdot\phi(Z_{n-1})}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$. Noting that $\frac{(Z_{n}-m\cdot\phi(Z_{n-1}))^{2}}{n\cdot\phi(Z_{n-1})}$ is non-negative, and that
    $\mathbbm{E}\Bigg{(}\frac{(Z_{n}-m\cdot\phi(Z_{n-1}))^{2}}{n\cdot\phi(Z_{n-1})}\Bigg{)}=\mathbbm{E}\Bigg{(}\frac{\text{Var}\big{(}Z_{n}\big{|}\mathcal{F}_{\phi_{n-1}})\big{)}}{n\cdot\phi(Z_{n-1})}\Bigg{)}=\frac{\sigma^{2}}{n}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
    this follows from Markov’s inequality.
(v) Strong consistency of $\boldsymbol{\bar{\beta}_{n}}$: Under the assumption that $\alpha$ is known, we introduce the estimator
$\tilde{\beta}_{n}:=\frac{(\phi(Z_{n})-\alpha\cdot Z_{n})^{2}}{Z_{n}},$
such that $\bar{\beta}_{n}=\frac{1}{n}\sum_{k=0}^{n-1}\tilde{\beta}_{k}$, and define
$U_{\phi_{n}}:=\tilde{\beta}_{n}-\beta,$
such that $\mathbbm{E}(U_{\phi_{n}}|\mathcal{F}_{\phi_{n-1}})=\mathbbm{E}\big{(}\mathbbm{E}(U_{\phi_{n}}|\mathcal{F}_{n})\big{|}\mathcal{F}_{\phi_{n-1}}\big{)}=0$ and $\mathbbm{E}|U_{\phi_{n}}|\leq 2\beta$, while
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}(U^{2}_{\phi_{n}}|\mathcal{F}_{\phi_{n-1}})=\sum_{n=1}^{\infty}\frac{1}{n^{2}}\Big{(}\mathbbm{E}(\tilde{\beta}^{2}_{n}|\mathcal{F}_{\phi_{n-1}})-\beta^{2}\Big{)}<\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}(\tilde{\beta}^{2}_{n}|\mathcal{F}_{\phi_{n-1}}).$
Under the assumption that there exists $d>0$ such that $\sup_{z\geq 1}\Big{\{}\frac{\mathbbm{E}(\phi(z)-\varepsilon(z))^{4}}{z^{2}}\Big{\}}\leq d$, we have that $\mathbbm{E}(\tilde{\beta}^{2}_{n}|\mathcal{F}_{\phi_{n-1}})\leq d$, so that
$\sum_{n=1}^{\infty}\frac{1}{n^{2}}\mathbbm{E}(U^{2}_{\phi_{n}}|\mathcal{F}_{\phi_{n-1}})<\sum_{n=1}^{\infty}\frac{d}{n^{2}}<\infty.$
Therefore, by Corollary 6.2, $\frac{1}{n}\sum_{k=1}^{n}U_{\phi_{k-1}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ and thus $\bar{\beta}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\beta$.
(vi) Weak consistency of $\boldsymbol{\hat{\beta}_{n}}$: We abandon the assumption that $\alpha$ is known. We can decompose
$\hat{\beta}_{n}=\bar{\beta}_{n}+\underbrace{\frac{2(\alpha-\tilde{\alpha}_{n-1})}{n}\sum_{k=0}^{n-1}(\phi(Z_{k})-\alpha Z_{k})}_{\text{I}}+\underbrace{\frac{(\alpha-\tilde{\alpha}_{n-1})^{2}}{n}\sum_{k=0}^{n-1}Z_{k},}_{\text{II}}$
where we have shown in (v) that $\bar{\beta}_{n}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}\beta$. It remains to show that I and II both converge in probability to zero.
@@
    We can decompose
    $\hskip-17.07182pt\frac{2(\alpha-\tilde{\alpha}_{n-1})}{n}\sum_{k=0}^{n-1}(\phi(Z_{k})-\alpha Z_{k})=-2\underbrace{\sqrt{\frac{\sum_{k=1}^{n-1}\phi(Z_{k-1})}{Z_{n-1}}}}_{\text{(a)}}\cdot\underbrace{\frac{\phi(Z_{n-1})-\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}}_{\text{(b)}}\cdot\underbrace{\frac{\sum_{k=0}^{n-1}(\phi(Z_{k})-\alpha Z_{k})}{n^{3/4}\sqrt{\sum_{k=1}^{n-1}\phi(Z_{k-1})}}}_{\text{(c)}},$
    and consider each of the components $(a)$, $(b)$ and $(c)$ in turn.
    1. (a)
      As a result of Lemma 6.5, for any $s$ such that $1<s<m\alpha$, we have that
      $\lim_{n\to\infty}\mathbbm{P}\Bigg{(}\sqrt{\frac{\sum_{k=1}^{n-1}\phi(Z_{k-1})}{\phi(Z_{n-2})}}\leq\sqrt{\frac{1}{1-s^{-1}}}\Bigg{)}=1.$
      Using Chebyshev’s inequality, for $r<m$,
      $\hskip-11.38092pt\mathbbm{P}(Z_{n-1}>r\cdot\phi(Z_{n-2})\,|\,\phi(Z_{n-2}))$ $\geq 1-\mathbbm{P}\big{(}|Z_{n-1}-m\cdot\phi(Z_{n-2})|\geq(m-r)\cdot\phi(Z_{n-2})\,\big{|}\,\phi(Z_{n-2})\big{)}$
      $\geq 1-\frac{\sigma^{2}}{(m-r)^{2}\cdot\phi(Z_{n-2})}$
      By conditioning on the event $\{Z_{n-1}>r\cdot\phi(Z_{n-2})\}$, we then find that
      $\lim_{n\to\infty}\mathbbm{P}\Bigg{(}\sqrt{\frac{\sum_{k=1}^{n-1}\phi(Z_{k-1})}{Z_{n-1}}}\leq\sqrt{\frac{r}{1-s^{-1}}}\Bigg{)}$
      $=\lim_{n\to\infty}\mathbbm{P}\Bigg{(}\sqrt{\frac{\sum_{k=1}^{n-1}\phi(Z_{k-1})}{Z_{n-1}}}\leq\sqrt{\frac{r}{1-s^{-1}}}\,\Bigg{|}\,Z_{n-1}>r\cdot\phi(Z_{n-2})\Bigg{)}\cdot\mathbbm{P}(Z_{n-1}>r\cdot\phi(Z_{n-2}))$
      $=1$
      on the event $\{Z_{n}\to\infty\}$. Hence we see that $\sqrt{\frac{\sum_{k=1}^{n-1}\phi(Z_{k-1})}{Z_{n-1}}}$ is asymptotically bounded by a constant in probability.
    2. (b)
      Given that
      $\mathbbm{E}\bigg{(}\frac{\phi(Z_{n-1})-\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\bigg{|}\mathcal{F}_{n-1}\bigg{)}=0\quad\text{and}\quad\text{Var}\bigg{(}\frac{\phi(Z_{n-1})-\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\bigg{|}\mathcal{F}_{n-1}\bigg{)}=\frac{\beta}{\sqrt{n}},$
      we see that
      $\mathbbm{E}\bigg{(}\frac{\phi(Z_{n-1})-\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\bigg{)}=0\quad\text{and}\quad\text{Var}\bigg{(}\frac{\phi(Z_{n-1})-\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\bigg{)}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
      so, as a consequence of Chebyshev’s inequality, it follows that
      $\frac{\phi(Z_{n-1})-\alpha Z_{n-1}}{n^{1/4}\cdot\sqrt{Z_{n-1}}}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0.$
    3. (c)
      Defining
      $U_{\phi_{n-1}}:=\phi(Z_{n-1})-\alpha Z_{n-1}\quad\text{and}\quad J_{n}:=n^{3/4}\sqrt{\sum_{k=1}^{n-1}\phi(Z_{k-1})},$
      we can see that $\{U_{\phi_{n-1}}\}_{n\in\mathbbm{N}_{1}}$ is a martingale difference sequence with respect to the filtration $\{\mathcal{F}_{\phi_{n}}\}_{n\in\mathbbm{N}_{0}}$, since $\mathbbm{E}(U_{\phi_{n-1}}|\mathcal{F}_{\phi_{n-2}})=0$ for all $n\in\mathbbm{N}_{1}$ and
      $\mathbbm{E}|U_{\phi_{n-1}}|$ $\leq\sqrt{\mathbbm{E}(U_{\phi_{n-1}}^{2})}$ (Hölder’s inequality)
      $=\sqrt{\mathbbm{E}(\text{Var}(\phi(Z_{n-1})|\mathcal{F}_{n-1}))}$
      $=\sqrt{\beta\cdot\mathbbm{E}(Z_{n-1})}$
      $<\infty,$
      and $\{J_{n}\}_{n\in\mathbbm{N}_{0}}$ is positive, non-decreasing, with $\lim_{n\to\infty}J_{n}=\infty$. Additionally,
      $\sum_{n=1}^{\infty}J_{n}^{-2}\mathbbm{E}(U_{\phi_{n-1}}^{2}|\mathcal{F}_{\phi_{n-2}})$ $=\sum_{n=1}^{\infty}\frac{\mathbbm{E}(\text{Var}(\phi(Z_{n-1})|\mathcal{F}_{n-1})|\mathcal{F}_{\phi_{n-2}})}{n^{3/2}\cdot\sum_{k=1}^{n-1}\phi(Z_{k-1})}$
      $=\sum_{n=1}^{\infty}\frac{m\beta\cdot\phi(Z_{n-2})}{n^{3/2}\cdot\sum_{k=1}^{n-1}\phi(Z_{k-1})}$
      $<\sum_{n=1}^{\infty}\frac{m\beta}{n^{3/2}}$
      $<\infty\;\;\text{a.s.},$
      so Theorem 6.1 applies, such that $J_{n}^{-1}\sum_{k=1}^{n}U_{\phi_{k-1}}=\frac{\sum_{k=0}^{n-1}(\phi(Z_{k})-\alpha Z_{k})}{n^{3/4}\sqrt{\sum_{k=1}^{n-1}\phi(Z_{k-1})}}\stackrel{{\scriptstyle a.s.}}{{\longrightarrow}}0$ as $n\to\infty$.
    Then, since $(a)$ is asymptotically bounded by a constant in probability, and $(b)$ and $(c)$ both converge to zero in probability, $\frac{2(\alpha-\tilde{\alpha}_{n-1})}{n}\sum_{k=0}^{n-1}(\phi(Z_{k})-\alpha Z_{k})\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$.
@@
    We can decompose
    $\frac{(\alpha-\tilde{\alpha}_{n-1})^{2}}{n}\sum_{k=0}^{n-1}Z_{k}=\frac{(\phi(Z_{n-1})-\alpha Z_{n-1})^{2}}{n\cdot Z_{n-1}}\cdot\sum_{k=1}^{n}\frac{Z_{k-1}}{Z_{n-1}},$
    and have already shown in Lemma 6.4 that $\sum_{k=1}^{n}\frac{Z_{k-1}}{Z_{n-1}}$ is asymptotically bounded in probability by a constant. Therefore the result $\frac{(\alpha-\tilde{\alpha}_{n-1})^{2}}{n}\sum_{k=0}^{n-1}Z_{k}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$ will follow if we show that $\frac{(\phi(Z_{n-1})-\alpha Z_{n-1})^{2}}{n\cdot Z_{n-1}}\stackrel{{\scriptstyle P}}{{\longrightarrow}}0$. Noting that $\frac{(\phi(Z_{n-1})-\alpha Z_{n-1})^{2}}{n\cdot Z_{n-1}}$ is non-negative, and that
    $\mathbbm{E}\Bigg{(}\frac{(\phi(Z_{n-1})-\alpha Z_{n-1})^{2}}{n\cdot Z_{n-1}}\Bigg{)}=\mathbbm{E}\Bigg{(}\frac{\text{Var}\big{(}\phi(Z_{n-1})\big{|}\mathcal{F}_{n-1})\big{)}}{n\cdot Z_{n-1}}\Bigg{)}=\frac{\beta}{n}\stackrel{{\scriptstyle n\to\infty}}{{\longrightarrow}}0,$
    this follows from Markov’s inequality.